### v1

# EEG-CLIP : Learning EEG representations from natural language descriptions

Abstract
Deep networks for electroencephalogram (EEG) decoding are currently often trained to only solve a specific task like pathology or gender decoding. A more general approach leveraging the medical reports of clinical EEG recordings is to learn mappings between medical reports and EEG recordings. This approach was pioneered in the computer vision domain matching images and their text captions and subsequently allowed to do successful zero-shot decoding using textual class prompts. In this work, we follow this approach and develop a contrastive learning framework EEG-CLIP that aligns EEG time series and their corresponding clinical text descriptions in a shared embedding space. We investigate its potential for versatile EEG decoding, assessing performance on a range of few-shot and zero-shot settings. Overall, results show that EEG-CLIP manages to nontrivially align text and EEG representations. Our work presents a promising approach to learn general EEG representations, which could enable easier analyses of diverse decoding questions through zero shot decoding or training task-specific models from fewer training examples. The code for reproducing our results is available at https://github.com/tidiane-camaret/EEGClip.
1 Introduction
Recent advances in machine learning have led to deep neural networks being commonly applied to electroencephalogram (EEG) data for a variety of decoding tasks. Roy et al. (2019) While deep learning models can achieve state-of-the-art performance on specialized EEG tasks, their learned representations often can only be used for one specific task. Most EEG analyses focus on training task-specific models for one type of classification or regression problem Heilmeyer et al., 2018. However, medical EEG recordings are often annotated with additional unstructured data in the form of free text reports written by neurologists and medical experts that can be exploited as a source of supervision.
In the computer vision domain, Contrastive Language–Image Pre-training Radford et al., 2021(CLIP) leverages text-image pairing to learn visual representations that effectively transfer across tasks. CLIP has demonstrated remarkable zero-shot generalization capabilities by learning to align images with natural language descriptions, enabling classification of previously unseen categories and adaptation to novel visual tasks without additional training.
Inspired by CLIP, we propose EEG-CLIP - a contrastive learning approach to align EEG time series data with corresponding clinical text descriptions in a shared embedding space. This work explores two central questions: (i) how textual reports can be incorporated into an EEG training pipeline, and (ii) to what extent this multimodal approach contributes to more general EEG representation learning.
We demonstrate EEG-CLIP’s potential for versatile EEG decoding through extensive evaluation on few-shot and zero-shot learning tasks. Our results show that EEG-CLIP achieves strong zero-shot classification performance and consistently outperforms previous transfer learning approaches and task-specific models when labeled data is scarce. This presents a promising direction for EEG analysis by enabling both zero-shot inference through natural language queries and more efficient training of specialized models with limited annotations.
2 Related Work
2.1 Deep-Learning based EEG decoding
2.2 CLIP : Alignment of data modalities
Contrastive self-supervised learning has recently emerged as a powerful approach for learning general visual representations. Models like CLIP are trained to align image and corresponding text embeddings by minimizing a contrastive loss :
where sim(.) is a measure of similarity. This objective brings matching image-text pairs closer and separates mismatched pairs in the learned embedding space.
CLIP is trained on a vast dataset of 400 million image-text pairs from diverse Internet sources with unstructured annotations. Through this natural language supervision, CLIP develops versatile image representations that achieve strong zero-shot inference on downstream tasks by querying the aligned embedding space.
The success of CLIP highlights the promise of contrastive learning approaches and leveraging readily available text data to learn transferable representations of other modalities like images. However, contrastive frameworks have not yet been explored for sequential data like EEG time series paired with text reports.
3 Experiments
3.1 Dataset
The Temple University Hospital EEG corpus (Obeid & Picone, 2016) contains over 25,000 EEG recordings collected from over 14,000 patients between 2002 and 2015. The large number of EEG recordings make this a valuable training dataset for deep learning models to learn to decode information such as pathology or age from the EEG and be able to generalize to unseen EEG recordings. The TUH Abnormal dataset (TUAB) is a demographically balanced subset with binary labels indicating pathological or nonpathological diagnosis of each recording. It is partitioned into training (1,387 normal and 1,398 abnormal files) and evaluation (150 normal and 130 abnormal files) sets. It contains a variety of pathological conditions.
Each recording contains additional labels : “age” (integer), “gender” (“M” or “F”), and “report” (string), a medical report written in natural language. The report is divided in 15 sections, listed in Table 1.
| Record section | Non-empty entries | Average word count |
| IMPRESSION | 2,971 | 16 |
| DESCRIPTION OF THE RECORD | 2,964 | 70 |
| CLINICAL HISTORY | 2,947 | 26 |
| MEDICATIONS | 2,893 | 4 |
| INTRODUCTION | 2,840 | 31 |
| CLINICAL CORRELATION | 2,698 | 31 |
| HEART RATE | 1,458 | 2 |
| FINDINGS | 887 | 16 |
| REASON FOR STUDY | 713 | 2 |
| TECHNICAL DIFFICULTIES | 684 | 3 |
| EVENTS | 569 | 8 |
| CONDITION OF THE RECORDING | 116 | 30 |
| PAST MEDICAL HISTORY | 19 | 8 |
| TYPE OF STUDY | 16 | 3 |
| ACTIVATION PROCEDURES | 9 | 3 |
3.2 EEG data preprocessing
We preprocess the EEG data, taking inspiration from the preprocessing steps in Schirrmeister et al. (2018) :
Select a subset of 21 electrodes present in all recordings.
Exclude the first minute of the recordings, and only use the first 2 minutes after that
CLIP the amplitude values to the range of ±800 V to reduce the effects of strong artifacts.
Resample the data to 100 Hz to further speed up the computation.
Divide by 30 to get closer to unit variance
3.3 Architecture and training details
The EEG-CLIP model is composed of two main components: an EEG encoder and a text encoder. These encoders are designed to process EEG recordings and medical reports respectively, as depicted in Figure 1.
For the EEG encoder, we use a convolutional neural network (CNN), Deep4 (Schirrmeister et al., 2018), whose architecture is optimized for the classification of EEG data. The Deep4 Network features four convolution-max-pooling blocks, using batch normalization and dropout, followed by a dense softmax classification layer. This enables the model to learn hierarchical spatial-temporal representations of the EEG signal. The output is flattened and passed to a fully-connected layer to derive a 128-dimensional embedding.
For the text encoder, we leverage pretrained text encoders based on the BERT architecture (Devlin et al., 2019). Such transformer-based models have shown state-of-the-art performance on a variety of natural language processing tasks. The advantage of these pretrained models is that they provide rich linguistic representations that can be effectively transferred to downstream tasks through finetuning.
The EEG and text embeddings are then fed into MLP projection heads, consisting of 3 fully-connected layers with ReLU activations. The final layer outputs a 64-dimensional projection of the embedding for contrastive learning. This architecture allows the model to learn alignments between EEG windows and corresponding medical report sentences in a shared embedding space. The contrastive loss enables the useful semantic features to be captured.
We train EEG-CLIP using the Adam optimizer with a learning rate of 5e-3 and weight decay of 5e-4. The model is trained for 20 epochs with a batch size of 64. We use the same training/testing split as in the TUAB dataset. Each recording is split in windows of length 1200, corresponding to a 12 second period, and with a stride of 519, which ensures all timesteps are predicted without any gap by our Deep4 model.
3.4 Evaluation methods
Unlike models trained for a specific downstream task, EEG-CLIP has to learn broadly useful representations that capture semantic relationships between EEG signals and text. As such, evaluation methods must aim to quantify the general quality and transferability of the learned representations.
Using the labels, meta information and medical reports provided in the TUAB dataset, we select 4 decoding tasks:
“pathological” : decode whether the recording was diagnosed as normal or pathological
“age”: decode whether the age of the patient is smaller or equal, or greater than 50
“gender”: decode the declared gender of the patient
“medication” : decode whether the medical report contains at least one of the 3 most common anticonvulsant medications ("keppra", "dilantin", "depakote")
We then design multiple methods to evaluate the model on, listed in the following.
3.4.1 Classification
In this method, we use the representations learned by the EEG encoder of EEG-CLIP as features for the 4 classification tasks. The encoder representations are kept frozen, while a classifier is trained and evaluated on the "train" and "eval" sections of TUAB. This enables us to see whether EEG-CLIP has learned to compress the discriminative information relevant for the classification tasks into its embeddings.
We choose to compare this approach against 2 baselines models with the same architecture as the EEG encoder:
The first baseline is a fully trainable model. This provides an upper bound on performance since the model can directly optimize for each task. Comparing to the trainable model reveals how much room there is for improvement over the fixed EEG-CLIP features.
The second baseline is a model trained on an unrelated task, which features are frozen while only the classification head is trainable. As these task-unrelated representations are not specialized for the actual decoding tasks, this serves as a lower bound on expected performance. The gap between the lower baseline and EEG-CLIP quantifies the benefits of our contrastive approach that can learn from all information contained in the medical reports.
By situating EEG-CLIP between these upper and lower bounds, we can better isolate the contributions of the learned representations themselves. Smaller gaps to the trainable model and larger gaps from the task-unrelated features indicate higher-quality multi-modal representations.
3.4.2 Zero-shot classification
We also perform zero-shot evaluation, using the embeddings of class-specific text prompts as class prototypes for the trained EEG-CLIP Model. For a given classification task, we define a typical prompt sentence for each class (see Table 2) and calculate the distance of an EEG recording to those sentences in the shared embedding space. This allows us to measure the classification performance of EEG-CLIP without any training on the classification task labels.
| Task | Prompt A | Prompt B |
| Pathological | “This is a normal recording” | “This is an abnormal recording” |
| Age | “The patient is under 50 years old” | “The patient is over 50 years old” |
| Gender | “The patient is male” | “The patient is female” |
| Medication | “No anti-epileptic drugs were prescribed to the patient” | “Anti-epileptic drugs were prescribed to the patient” |
3.4.3 Classification in a low-data regime
To further evaluate the generalization capability of the learned representations, we assess few-shot performance by training the classifier on a small subset, held-out from the training of EEG-CLIP. The limited labeled data setting reflects realistic clinical scenarios where large labeled datasets are difficult to acquire. New clinical applications often only have access to small patient datasets. As such, assessing few-shot transfer is important for demonstrating clinical utility and feasibility.
4 Results
4.1 Evaluation of the learned representations
In this section, we present results evaluating the learned representations from EEG-CLIP across a diverse set of experiments. As a reminder, our evaluation methodology consisted of classification tasks using the full TUAB dataset, zero-shot classification using text prompts, few-shot classification on a held-out dataset, and a recording retrieval task.
4.1.1 Classification
As shown in 3, on three of the four tasks, EEG-CLIP with a simple logistic regression classifier achieved strong performance, with balanced accuracy scores of 0.826 for pathological status, 0.687 for gender, and 0.713 for age. This indicates the representations capture meaningful signal related to these key attributes.
With a 3-layer MLP classifier head, performance improved further on all tasks, reaching 0.847, 0.702, and 0.747 respectively. The MLP can better exploit the relationships in the embedding space.
As expected, the task-specific models achieve the top scores, as they are optimized end-to-end directly on the evaluation data distribution and labels. However, EEG-CLIP comes reasonably close, especially with the MLP head, demonstrating the generalizability of the representations.
Compared to irrelevant task pretraining, EEG-CLIP substantially outperforms models pretrained on inconsistent targets like age or pathology. This confirms the importance of learning from the information contained in the medical reports.
| Task | EEG-CLIP | Task-specific | Irrelevant task | |
| LogReg | MLP | |||
| Pathological | 0.826 | 0.847 | 0.851 | 0.741 (age) |
| Gender | 0.687 | 0.702 | 0.752 | 0.667 (pathological) |
| Age | 0.713 | 0.747 | 0.786 | 0.685 (pathological) |
| Medication | 0.633 | 0.615 | 0.685 | 0.573 (pathological) |
4.1.2 Zero-shot classification
For zero-shot classification, we evaluate EEG-CLIP’s ability to classify EEG recordings without any task-specific training. We create text prompts that describe each class (see Table 2) and compute similarities between EEG embeddings and these text prompts in the shared embedding space. As shown in Table 4, EEG-CLIP achieves remarkable zero-shot performance on the pathological task (0.755), demonstrating strong alignment between EEG signals and their textual descriptions. Performance on age classification (0.642) is also substantially above chance, while gender (0.567) and medication (0.532) show more modest improvements. These results are particularly encouraging as they represent classification without any labeled training data, relying solely on the semantic alignment learned during contrastive training. The strong pathology detection performance suggests that diagnostic language in the medical reports is effectively aligned with corresponding neurological patterns in the EEG signals. This zero-shot capability could be especially valuable in clinical settings where labeled data for new tasks is scarce or unavailable.
| Task | Accuracy |
| Pathological | 0.755 |
| Age | 0.642 |
| Gender | 0.567 |
| Medication | 0.532 |
4.1.3 Classification in a low-data regime
| Task | EEG-CLIP + MLP | Task-specific | Irrelevant task |
| Pathological | 0.710 | 0.781 | 0.531 (age) |
| Gender | 0.550 | 0.648 | 0.512 (pathological) |
| Age | 0.712 | 0.621 | 0.631 (pathological) |
| Medication | 0.551 | 0.575 | 0.598 (pathological) |
On the pathological task, EEG-CLIP achieves 0.710 balanced accuracy on the held-out set. This approaches the 0.781 performance of a model trained from scratch with the same limited data. For age classification, EEG-CLIP even outperforms the specialized model. The medication task proves most challenging in the few-shot setting. However, all models struggle to exceed 0.6 accuracy, suggesting intrinsic difficulty of the binary prediction from small samples.
Critically, EEG-CLIP substantially outperforms models pretrained on irrelevant tasks across all but one experiment. This demonstrates the concrete value of pretraining on aligned data, even when fine-tuning data is scarce.
Taken together, these quantitative results provide strong evidence for the quality and transferability of the multi-modal representations learned by EEG-CLIP. Performance across the range of evaluation paradigms demonstrates that EEG-CLIP successfully encodes general semantic relationships between EEG and text. This enables the model to generalize to new tasks and datasets without task-specific fine-tuning.
4.2 Impact of the report sections on the representations
In order to analyze the relative impact of different report categories on learning effective EEG representations, we conduct ablation experiments training separate models with single category inclusions. Specifically, we feed impressions-only, descriptions-only, or history-only texts as supervision for the contrastive embedding process across separate model instances. Keeping the methodology constant, these category-specific versions reveal how report sections influences model performance.
As expected, training with all report categories leads to the overall best model performance across tasks. However, our ablation study reveals some unexpected category-specific advantages indicating useful correlations in the data. For example, a heart rate-only model variant achieves higher accuracy in decoding patient gender. Though counterintuitive, this result suggests heart rate descriptors alone better capture differences in cardiac patterns between males and females. Similarly, a technical difficulties model focusing on recording quality notes gains an edge in anomaly detection - learning to recognize atypical artifacts.
Overall, while impression, description, and history make the largest individual contributions, every report section provides some specialization. Our experiments uncover these subtle inter-relationships across text components and EEG decoding objectives. Certain specialized descriptors, when isolated, help the model tune into narrow physiological signals that happen to be particularly predictive for certain tasks. Hence, combining categories is ideal, but analysis also shows how targeting select report data can benefit specific use cases when limited modalities are available.
4.3 Study on parameter importance
4.3.1 Pre-training and training of the text encoder
We experimented with several variants including bert-base-uncased, a general domain model trained on Wikipedia and BookCorpus; ClinicalBERT, a model finetuned on clinical notes; BioGPT-Large-PubMedQA, tailored for biomedical text; and BGE-Large, a model trained on scientific papers and designed for generation tasks.
4.3.2 Projected embedding dimension
Additionally, we analyze EEG-CLIP model performance across varied hidden dimensionality sizes for the jointly learned EEG and text embeddings. Counter to typical representation learning trends, we find higher decoding accuracy with smaller shared embedding spaces between 32-128 dimensions rather than larger 256 or 512 sizes. Compressing the modalities into more compact unified vectors likely distills the patterns into their most essential latent characteristics necessary for generalization. Meanwhile, overparameterization could enable overfitting particular artifacts of the training distribution. Further, smaller spaces may inherently force more direct alignment between descriptive concepts and neurological patterns. By tuning towards optimal lower-complexity manifolds, EEG-CLIP better transfers key aspects in the aligned embeddings while filtering out unnecessary signals.
5 Discussion
Our experiments demonstrate that EEG-CLIP successfully learns to align EEG recordings and their clinical text descriptions in a shared embedding space. This approach shows promise for developing more versatile and generalizable EEG representations that can transfer across multiple decoding tasks.
The model achieves strong performance on standard classification tasks (balanced accuracies of 0.847 for pathology, 0.702 for gender, and 0.747 for age) when using an MLP classifier head. Most notably, EEG-CLIP demonstrates zero-shot classification capabilities, achieving 0.755 balanced accuracy on pathology detection using only natural language prompts without any task-specific training. In low-data regimes, our approach shows substantial benefits over models trained from scratch or pretrained on irrelevant tasks, suggesting efficient capture of generalizable features.
Our ablation studies reveal that while using all report sections yields the best overall performance, specific sections provide distinct advantages for certain tasks. Interestingly, we found that smaller embedding dimensions (32-128) outperformed larger ones, contrary to common intuition in representation learning. This suggests that compressing information into a more compact shared embedding space may better distill essential cross-modal patterns.
Despite these promising results, several limitations warrant discussion. Our experiments were limited to the TUAB dataset, which may not capture the full diversity of EEG patterns and neurological conditions. Clinical reports in the dataset vary in quality, detail, and structure, potentially limiting the model’s ability to learn precise alignments. Additionally, our current approach treats entire EEG recordings and their corresponding reports as aligned pairs, whereas more fine-grained temporal alignment could further improve performance.
Future work could explore methods for aligning specific EEG segments with relevant sentences in clinical reports, scaling to larger and more diverse EEG datasets, incorporating additional data modalities, and developing methods to interpret the learned representations in terms of clinically meaningful EEG patterns. The development of foundation models for biosignals, analogous to large language models or vision transformers, represents an exciting direction for this research.
In conclusion, EEG-CLIP demonstrates the feasibility of contrastive learning between EEG signals and natural language descriptions for developing more general and transferable EEG representations. While this represents just an initial exploration, the approach opens up new possibilities for leveraging unstructured clinical text to enhance deep learning models for EEG analysis, potentially leading to more flexible, efficient, and powerful tools for neurological assessment and research.

### v2

@@
 Abstract
-Deep networks for electroencephalogram (EEG) decoding are currently often trained to only solve a specific task like pathology or gender decoding.
-A more general approach leveraging the medical reports of clinical EEG recordings is to learn mappings between medical reports and EEG recordings.
+Deep networks for electroencephalogram (EEG) decoding are often only trained to solve one specific task, such as pathology or age decoding.
+A more general task-agnostic approach is to train deep networks to match a (clinical) EEG recording to its corresponding textual medical report and vice versa.
 This approach was pioneered in the computer vision domain matching images and their text captions and subsequently allowed to do successful zero-shot decoding using textual class prompts.
-In this work, we follow this approach and develop a contrastive learning framework EEG-CLIP that aligns EEG time series and their corresponding clinical text descriptions in a shared embedding space.
-We investigate its potential for versatile EEG decoding, assessing performance on a range of few-shot and zero-shot settings.
-Overall, results show that EEG-CLIP manages to nontrivially align text and EEG representations.
-Our work presents a promising approach to learn general EEG representations, which could enable easier analyses of diverse decoding questions through zero shot decoding or training task-specific models from fewer training examples.
+In this work, we follow this approach and develop a contrastive learning framework, EEG-CLIP, that aligns the EEG time series and the descriptions of the corresponding clinical text in a shared embedding space.
+We investigated its potential for versatile EEG decoding, evaluating performance in a range of few-shot and zero-shot settings.
+Overall, we show that EEG-CLIP manages to non-trivially align text and EEG representations.
+Our work presents a promising approach to learn general EEG representations, which could enable easier analyses of diverse decoding questions through zero-shot decoding or training task-specific models from fewer training examples.
 The code for reproducing our results is available at https://github.com/tidiane-camaret/EEGClip.
+Keywords: Electroencephalogram (EEG), Contrastive Learning, Multimodal Representation, Zero-shot Classification, Clinical Text Processing, Neural Time Series, Transfer Learning
 1 Introduction
-Recent advances in machine learning have led to deep neural networks being commonly applied to electroencephalogram (EEG) data for a variety of decoding tasks.
-Roy et al.
-(2019) While deep learning models can achieve state-of-the-art performance on specialized EEG tasks, their learned representations often can only be used for one specific task.
-Most EEG analyses focus on training task-specific models for one type of classification or regression problem Heilmeyer et al., 2018.
-However, medical EEG recordings are often annotated with additional unstructured data in the form of free text reports written by neurologists and medical experts that can be exploited as a source of supervision.
-In the computer vision domain, Contrastive Language–Image Pre-training Radford et al., 2021(CLIP) leverages text-image pairing to learn visual representations that effectively transfer across tasks.
+Recent advances in machine learning have led to deep neural networks being commonly applied to electroencephalogram (EEG) data for a variety of decoding tasks (Roy et al., 2019).
+While deep learning models can achieve state-of-the-art performance on specialized EEG tasks, their learned representations can often only be used for one specific task.
+Most EEG analyses focus on training task-specific models for one type of classification or regression problem (Heilmeyer et al., 2018).
+However, medical EEG recordings are often accompanied by rich unstructured annotations in the form of free text reports written by neurologists and medical experts—a potentially valuable source of supervision that remains largely untapped.
+In the computer vision domain, Contrastive Language–Image Pre-training or CLIP (Radford et al., 2021) leverages text-image pairing to learn visual representations that effectively transfer across tasks.
 CLIP has demonstrated remarkable zero-shot generalization capabilities by learning to align images with natural language descriptions, enabling classification of previously unseen categories and adaptation to novel visual tasks without additional training.
-Inspired by CLIP, we propose EEG-CLIP - a contrastive learning approach to align EEG time series data with corresponding clinical text descriptions in a shared embedding space.
-This work explores two central questions: (i) how textual reports can be incorporated into an EEG training pipeline, and (ii) to what extent this multimodal approach contributes to more general EEG representation learning.
+Inspired by CLIP, we propose EEG-CLIP : a contrastive learning approach to align EEG time-series data with corresponding clinical text descriptions in a shared embedding space.
+This work explores two central questions: (i) how clinical text reports can be effectively incorporated into EEG representation learning, and (ii) whether this multimodal approach enables more generalizable representations that transfer across diverse EEG decoding tasks.
 We demonstrate EEG-CLIP’s potential for versatile EEG decoding through extensive evaluation on few-shot and zero-shot learning tasks.
-Our results show that EEG-CLIP achieves strong zero-shot classification performance and consistently outperforms previous transfer learning approaches and task-specific models when labeled data is scarce.
-This presents a promising direction for EEG analysis by enabling both zero-shot inference through natural language queries and more efficient training of specialized models with limited annotations.
+Our results show that EEG-CLIP achieves strong zero-shot classification performance and consistently outperforms previous transfer learning approaches and task-specific models when labeled data are scarce.
+This presents a promising direction for EEG analysis by enabling zero-shot inference through natural language queries and more efficient training of specialized models with limited annotations.
+Remark Recently and after the completion of the study presented in this manuscript, (Gijsen & Ritter, 2024) also proposed EEG-language models that align EEG data with clinical reports for pathology detection.
+Their work explores multiple alignment strategies, including a multiple instance learning extension for flexible matching between EEG segments and text portions.
+Their approach is primarily focused on pathology detection and classification of epileptiform activity, while our EEG-CLIP study examines model performance on diverse decoding objectives including age, gender, and medication prediction, providing further insights into the versatility of language-supervised EEG representations.
 2 Related Work
 2.1 Deep-Learning based EEG decoding
-2.2 CLIP : Alignment of data modalities
-Contrastive self-supervised learning has recently emerged as a powerful approach for learning general visual representations.
-Models like CLIP are trained to align image and corresponding text embeddings by minimizing a contrastive loss :
-where sim(.) is a measure of similarity. This objective brings matching image-text pairs closer and separates mismatched pairs in the learned embedding space.
-CLIP is trained on a vast dataset of 400 million image-text pairs from diverse Internet sources with unstructured annotations.
-Through this natural language supervision, CLIP develops versatile image representations that achieve strong zero-shot inference on downstream tasks by querying the aligned embedding space.
-The success of CLIP highlights the promise of contrastive learning approaches and leveraging readily available text data to learn transferable representations of other modalities like images.
-However, contrastive frameworks have not yet been explored for sequential data like EEG time series paired with text reports.
-3 Experiments
+Deep learning has revolutionized EEG analysis by enabling end-to-end decoding directly from raw signals without hand-crafted features.
+Convolutional neural networks (CNNs) have shown particular promise, with recent advances like batch normalization and exponential linear units boosting performance to match or exceed traditional methods like filter bank common spatial patterns (FBCSP) (Schirrmeister et al., 2017) These architectures automatically learn hierarchical representations that capture relevant spectral and spatial patterns in EEG data.
+Various neural network architectures have been applied to EEG tasks, from shallow CNNs for efficient processing to recurrent networks for capturing temporal dependencies.
+Recent comparative studies have demonstrated that specialized deep learning models can outperform traditional approaches on standard benchmarks such as BCI Competition datasets.
+Beyond classification, newer approaches like EEG-to-text decoding leverage advanced neural architectures combined with probabilistic modeling to translate neural activity into human-readable text (Lévy et al., 2025), expanding the potential applications for EEG-based interfaces.
+Multimodal approaches have also shown promise, such as Khan et al.
+(2018) who combined EEG with fNIRS to capture complementary neurophysiological signals.
+While they used traditional signal processing rather than deep learning, their work highlights the value of integrating EEG with additional information sources, a principle that motivates our EEG-text alignment approach.
+While task-specific models dominate current approaches, our work explores a more general representation learning framework that leverages the rich information in clinical text reports to develop versatile EEG embeddings useful across multiple decoding tasks.
+2.2 Contrastive Learning for Multimodal Alignment
+Self-supervised contrastive learning has recently emerged as a powerful approach to learning general visual representations.
+Models like CLIP are trained to align the image embeddings and the corresponding text embeddings by minimizing contrast loss :
+where is a measure of similarity. This objective brings the matching image-text pairs closer and separates the mismatched pairs in the learned embedding space.
+CLIP was trained on a large dataset of 400 million image-text pairs from diverse Internet sources with unstructured annotations.
+Through this natural language supervision, CLIP developed versatile image representations that achieve strong zero-shot inference on downstream tasks by querying the aligned embedding space.
+The success of CLIP highlights the promise of contrastive learning approaches and the use of readily available text data to learn transferable representations of other modalities.
+3 Methods
 3.1 Dataset
@@
 3.2 EEG data preprocessing
-We preprocess the EEG data, taking inspiration from the preprocessing steps in Schirrmeister et al. (2018) :
+We preprocess the EEG data, following the preprocessing steps from Schirrmeister et al. (2017) :
 Select a subset of 21 electrodes present in all recordings.
@@
 These encoders are designed to process EEG recordings and medical reports respectively, as depicted in Figure 1.
-For the EEG encoder, we use a convolutional neural network (CNN), Deep4 (Schirrmeister et al., 2018), whose architecture is optimized for the classification of EEG data.
+For the EEG encoder, we use a convolutional neural network (CNN), Deep4 (Schirrmeister et al., 2017), whose architecture is optimized for the classification of EEG data.
 The Deep4 Network features four convolution-max-pooling blocks, using batch normalization and dropout, followed by a dense softmax classification layer.
@@
 The contrastive loss enables the useful semantic features to be captured.
-We train EEG-CLIP using the Adam optimizer with a learning rate of 5e-3 and weight decay of 5e-4.
+We train EEG-CLIP using the Adam optimizer with a learning rate of and weight decay of .
 The model is trained for 20 epochs with a batch size of 64.
@@
 As such, evaluation methods must aim to quantify the general quality and transferability of the learned representations.
-Using the labels, meta information and medical reports provided in the TUAB dataset, we select 4 decoding tasks:
+Using the labels and medical reports provided in the TUAB dataset, we select 4 decoding tasks:
 “Pathological” : decode whether the recording was diagnosed as normal or pathological
@@
 “Gender”: decode the declared gender of the patient
-“medication” : decode whether the medical report contains at least one of the 3 most common anticonvulsant medications ("keppra", "dilantin", "depakote")
+“Medication” : decode whether the medical report contains at least one of the 3 most common anticonvulsant medications ("keppra", "dilantin" and "depakote")
 We then design multiple methods to evaluate the model on, listed in the following.
 3.4.1 Classification
-In this method, we use the representations learned by the EEG encoder of EEG-CLIP as features for the 4 classification tasks.
-The encoder representations are kept frozen, while a classifier is trained and evaluated on the "train" and "eval" sections of TUAB.
-This enables us to see whether EEG-CLIP has learned to compress the discriminative information relevant for the classification tasks into its embeddings.
-We choose to compare this approach against 2 baselines models with the same architecture as the EEG encoder:
-The first baseline is a fully trainable model.
-This provides an upper bound on performance since the model can directly optimize for each task.
-Comparing to the trainable model reveals how much room there is for improvement over the fixed EEG-CLIP features.
-The second baseline is a model trained on an unrelated task, which features are frozen while only the classification head is trainable.
-As these task-unrelated representations are not specialized for the actual decoding tasks, this serves as a lower bound on expected performance.
-The gap between the lower baseline and EEG-CLIP quantifies the benefits of our contrastive approach that can learn from all information contained in the medical reports.
-By situating EEG-CLIP between these upper and lower bounds, we can better isolate the contributions of the learned representations themselves.
-Smaller gaps to the trainable model and larger gaps from the task-unrelated features indicate higher-quality multi-modal representations.
+We compare EEG-CLIP against two baseline models to contextualize its performance:
+Task-specific model (upper bound): A Deep4 CNN trained end-to-end from random initialization directly on each target task.
+This provides an upper bound since the entire model can optimize specifically for the task.
+Alternative task transfer model (lower bound): A Deep4 CNN first trained from random initialization on an unrelated task (e.g., age classification), then transferred to the target task (e.g., pathology detection) by freezing the encoder and training only a new classification head.
+This tests whether simple transfer learning from any EEG task provides useful features.
+Both baselines use the same architecture as EEG-CLIP’s EEG encoder but differ in their training approach: EEG-CLIP uses contrastive learning with text supervision, while baselines use standard supervised learning with task labels.
+Figure 2 illustrates the different training strategies.
 3.4.2 Zero-shot classification
@@
 3.4.3 Classification in a low-data regime
-To further evaluate the generalization capability of the learned representations, we assess few-shot performance by training the classifier on a small subset, held-out from the training of EEG-CLIP.
-The limited labeled data setting reflects realistic clinical scenarios where large labeled datasets are difficult to acquire.
-New clinical applications often only have access to small patient datasets.
-As such, assessing few-shot transfer is important for demonstrating clinical utility and feasibility.
+To further evaluate the generalization capability of the learned representations, we assess few-shot performance by training classifiers on varying fractions of a small labeled dataset. Specifically:
+We hold out 20% of the TUAB training set exclusively for few-shot training (never seen during EEG-CLIP’s contrastive training)
+From this 20% held-out set, we create subsets of sizes: 1/2, 1/5, 1/10, 1/20, and 1/50
+All models (EEG-CLIP, task-specific, and alternative task) are trained on these identical data subsets.
+The key difference is that EEG-CLIP uses representations learned from contrastive training on the separate 60% split, while the task-specific model trains from scratch.
+This ensures a fair comparison where all models have access to the same limited labeled data, isolating the benefit of pre-training.
+Figure 3 illustrates the data partitioning strategy.
 4 Results
@@
 In this section, we present results evaluating the learned representations from EEG-CLIP across a diverse set of experiments.
-As a reminder, our evaluation methodology consisted of classification tasks using the full TUAB dataset, zero-shot classification using text prompts, few-shot classification on a held-out dataset, and a recording retrieval task.
-4.1.1 Classification
-As shown in 3, on three of the four tasks, EEG-CLIP with a simple logistic regression classifier achieved strong performance, with balanced accuracy scores of 0.826 for pathological status, 0.687 for gender, and 0.713 for age.
-This indicates the representations capture meaningful signal related to these key attributes.
-With a 3-layer MLP classifier head, performance improved further on all tasks, reaching 0.847, 0.702, and 0.747 respectively. The MLP can better exploit the relationships in the embedding space.
-As expected, the task-specific models achieve the top scores, as they are optimized end-to-end directly on the evaluation data distribution and labels.
-However, EEG-CLIP comes reasonably close, especially with the MLP head, demonstrating the generalizability of the representations.
-Compared to irrelevant task pretraining, EEG-CLIP substantially outperforms models pretrained on inconsistent targets like age or pathology.
-This confirms the importance of learning from the information contained in the medical reports.
-| Task | EEG-CLIP | Task-specific | Irrelevant task | |
+As a reminder, our evaluation methodology consisted of classification tasks using the full TUAB dataset, zero-shot classification using text prompts, and few-shot classification on a held-out dataset.
+4.1.1 Classification performance
+Table 3 shows EEG-CLIP’s classification performance across four tasks.
+With logistic regression, EEG-CLIP achieves balanced accuracies of 0.826 for pathological status, 0.713 for age, and 0.687 for gender.
+A 3-layer MLP classifier further improves results to 0.847, 0.747, and 0.702 respectively, indicating non-linear relationships in the embedding space.
+The performance gap between EEG-CLIP+MLP and task-specific models remains small (0.004 for pathological, 0.039 for age, 0.050 for gender) despite the latter’s end-to-end optimization advantage.
+Most importantly, EEG-CLIP consistently outperforms alternative task pretraining by 10.6% for pathological, 6.2% for age, and 3.5% for gender classification.
+These quantitative results demonstrate that text-EEG contrastive learning produces more transferable representations than single-task supervised learning.
+| Task | EEG-CLIP | Task-specific | Alternative task | |
 | LogReg | MLP | |||
 | Pathological | 0.826 | 0.847 | 0.851 | 0.741 (age) |
+| Age | 0.713 | 0.747 | 0.786 | 0.685 (pathological) |
 | Gender | 0.687 | 0.702 | 0.752 | 0.667 (pathological) |
-| Age | 0.713 | 0.747 | 0.786 | 0.685 (pathological) |
 | Medication | 0.633 | 0.615 | 0.685 | 0.573 (pathological) |
-4.1.2 Zero-shot classification
+4.1.2 Zero-shot classification performance
 For zero-shot classification, we evaluate EEG-CLIP’s ability to classify EEG recordings without any task-specific training.
-We create text prompts that describe each class (see Table 2) and compute similarities between EEG embeddings and these text prompts in the shared embedding space.
+We compute similarities between EEG embeddings and text prompts (see Table 2) in the shared embedding space.
 As shown in Table 4, EEG-CLIP achieves remarkable zero-shot performance on the pathological task (0.755), demonstrating strong alignment between EEG signals and their textual descriptions.
-Performance on age classification (0.642) is also substantially above chance, while gender (0.567) and medication (0.532) show more modest improvements.
+Performance on age classification (0.642) is also substantially above chance, while gender (0.567) and medication (0.532) show more modest scores.
 These results are particularly encouraging as they represent classification without any labeled training data, relying solely on the semantic alignment learned during contrastive training.
@@
 | Medication | 0.532 |
-4.1.3 Classification in a low-data regime
-| Task | EEG-CLIP + MLP | Task-specific | Irrelevant task |
-| Pathological | 0.710 | 0.781 | 0.531 (age) |
-| Gender | 0.550 | 0.648 | 0.512 (pathological) |
-| Age | 0.712 | 0.621 | 0.631 (pathological) |
-| Medication | 0.551 | 0.575 | 0.598 (pathological) |
+4.1.3 Few-shot classification performance
 On the pathological task, EEG-CLIP achieves 0.710 balanced accuracy on the held-out set.
@@
 However, all models struggle to exceed 0.6 accuracy, suggesting intrinsic difficulty of the binary prediction from small samples.
-Critically, EEG-CLIP substantially outperforms models pretrained on irrelevant tasks across all but one experiment.
+The detailed results are presented in Table 5.
+| Task | EEG-CLIP + MLP | Task-specific | Alternative task |
+| Pathological | 0.710 | 0.781 | 0.531 (age) |
+| Age | 0.712 | 0.621 | 0.631 (pathological) |
+| Gender | 0.550 | 0.648 | 0.512 (pathological) |
+| Medication | 0.551 | 0.575 | 0.598 (pathological) |
+Critically, EEG-CLIP substantially outperforms models pretrained on alternative tasks across all but one experiment.
 This demonstrates the concrete value of pretraining on aligned data, even when fine-tuning data is scarce.
+As shown in Figure 4, EEG-CLIP (green lines) maintains relatively stable performance across increasingly smaller fractions of the training set, from down to of the original dataset.
+For pathology detection (top left), EEG-CLIP maintains strong performance even with minimal data (), outperforming both baselines as data becomes extremely scarce.
+Age classification (top right) shows EEG-CLIP consistently outperforming other approaches across all data regimes.
+For gender and medication tasks (bottom panels), all models show performance degradation with reduced data, but EEG-CLIP demonstrates greater robustness to extreme data reductions ().
 Taken together, these quantitative results provide strong evidence for the quality and transferability of the multi-modal representations learned by EEG-CLIP.
@@
 This enables the model to generalize to new tasks and datasets without task-specific fine-tuning.
+The recent ELM-MIL approach by Gijsen & Ritter (2024) achieves superior performance (87.11% balanced accuracy) through Multiple Instance Learning extensions that address fine-grained EEG-text alignment.
+While our approach achieves 84.7%, both works demonstrate that multimodal language supervision significantly outperforms EEG-only baselines, validating this research direction.
 4.2 Impact of the report sections on the representations
-In order to analyze the relative impact of different report categories on learning effective EEG representations, we conduct ablation experiments training separate models with single category inclusions.
-Specifically, we feed impressions-only, descriptions-only, or history-only texts as supervision for the contrastive embedding process across separate model instances.
-Keeping the methodology constant, these category-specific versions reveal how report sections influences model performance.
-As expected, training with all report categories leads to the overall best model performance across tasks.
-However, our ablation study reveals some unexpected category-specific advantages indicating useful correlations in the data.
-For example, a heart rate-only model variant achieves higher accuracy in decoding patient gender.
-Though counterintuitive, this result suggests heart rate descriptors alone better capture differences in cardiac patterns between males and females.
-Similarly, a technical difficulties model focusing on recording quality notes gains an edge in anomaly detection - learning to recognize atypical artifacts.
-Overall, while impression, description, and history make the largest individual contributions, every report section provides some specialization.
-Our experiments uncover these subtle inter-relationships across text components and EEG decoding objectives.
-Certain specialized descriptors, when isolated, help the model tune into narrow physiological signals that happen to be particularly predictive for certain tasks.
-Hence, combining categories is ideal, but analysis also shows how targeting select report data can benefit specific use cases when limited modalities are available.
+To analyze how different report sections influence EEG representation learning, we conducted systematic ablation experiments by training separate models with single section inputs.
+Each model utilized identical methodology except for the textual input, which was restricted to specific report sections (impression, description, history, etc.).
+We also experimented with randomly sampling sub-strings from each section during training, but this approach led to decreased performance compared to using complete sections.
+As shown in Figure 5, while using all report sections yielded the best overall performance across tasks, certain section-specific models demonstrated unexpected strengths.
+Notably, the heart rate section model achieved superior accuracy in gender classification despite its brevity (average 2 words per report), suggesting cardiac pattern descriptions capture gender-specific physiological differences.
+Similarly, a model trained exclusively on technical difficulties notes showed enhanced sensitivity to pathological recordings, likely by learning to associate recording artifacts with abnormal brain activity.
+The three primary sections (impression, description, and clinical history) provided the strongest individual contributions to performance, aligning with their higher word counts and prevalence across the dataset (Table 1).
+However, even sections with limited representation, such as medication lists, contributed unique predictive signals for specific tasks.
+These findings reveal how specialized clinical descriptions, even when isolated, can help models detect task-relevant physiological patterns in EEG data.
+While combining all sections remains optimal for general-purpose representations, our analysis demonstrates the potential value of targeting specific report sections when developing specialized decoders or when working with incomplete clinical documentation.
 4.3 Study on parameter importance
-4.3.1 Pre-training and training of the text encoder
-We experimented with several variants including bert-base-uncased, a general domain model trained on Wikipedia and BookCorpus; ClinicalBERT, a model finetuned on clinical notes; BioGPT-Large-PubMedQA, tailored for biomedical text; and BGE-Large, a model trained on scientific papers and designed for generation tasks.
+4.3.1 Pre-training of the text encoder
+We experimented with several variants of pre-trained language models as text encoders, available publicly on the Hugging Face Hub (Wolf et al., 2020), as shown in Figure 6.
+These included BERT-base-uncased (Devlin et al., 2018), a general domain model trained on Wikipedia and BookCorpus; ClinicalBERT (Huang et al., 2019), a model finetuned on clinical notes; BioGPT-Large-PubMedQA (Luo et al., 2022), tailored for biomedical text; and BGE-Large (Xiao et al., 2023), a model trained on scientific papers and designed for generation tasks.
+ClinicalBERT demonstrated the strongest overall performance across tasks, highlighting the advantage of domain-specific pre-training for clinical text processing.
+Interestingly, the general-purpose BERT-base-uncased maintained competitive performance on pathology detection despite lacking medical specialization.
+BGE-Large showed particular strength in pathology and age classification tasks, while BioGPT-Large-PubMedQA consistently underperformed across all evaluations.
+These results emphasize how encoder architecture and pre-training domain significantly impact the quality of cross-modal representations in our EEG-text alignment framework.
+The learning rate ratio between the text encoder and EEG encoder also proved critical, as shown in Figure 7.
+Optimal performance was achieved when the text encoder was fine-tuned at times the learning rate of the EEG encoder, balancing adaptation of pre-trained linguistic knowledge while preserving domain-specific understanding.
 4.3.2 Projected embedding dimension
-Additionally, we analyze EEG-CLIP model performance across varied hidden dimensionality sizes for the jointly learned EEG and text embeddings.
-Counter to typical representation learning trends, we find higher decoding accuracy with smaller shared embedding spaces between 32-128 dimensions rather than larger 256 or 512 sizes.
-Compressing the modalities into more compact unified vectors likely distills the patterns into their most essential latent characteristics necessary for generalization.
-Meanwhile, overparameterization could enable overfitting particular artifacts of the training distribution.
-Further, smaller spaces may inherently force more direct alignment between descriptive concepts and neurological patterns.
-By tuning towards optimal lower-complexity manifolds, EEG-CLIP better transfers key aspects in the aligned embeddings while filtering out unnecessary signals.
+Additionally, we analyzed EEG-CLIP model performance across varied hidden dimensionality sizes for the jointly learned EEG and text embeddings, as illustrated in Figure 8.
+Counter to typical representation learning trends, we found higher decoding accuracy with smaller shared embedding spaces between 32-128 dimensions rather than larger 256 or 512 sizes.
+A t-SNE visualization of the 64-dimensional embeddings in Figure 9 reveals clear clustering by pathological status, demonstrating effective semantic organization of the learned representation space.
+This counterintuitive finding suggests that compressing both modalities into compact unified vectors distills patterns into their most essential characteristics necessary for generalization, while avoiding overfitting to training distribution artifacts that may occur in higher-dimensional spaces.
+The constrained dimensionality may also enforce more direct alignment between descriptive clinical concepts and underlying neurological patterns.
+These results indicate that EEG-CLIP benefits from lower-complexity manifolds that capture key cross-modal correspondences while filtering out extraneous signals.
+4.4 Gradient Analysis for Model Interpretability
+In order to provide insights into EEG-CLIP’s decision-making process, we performed gradient-based analysis to visualize which EEG regions contribute most strongly to embeddings aligned with specific textual concepts.
+We computed gradients of the cosine similarity between EEG embeddings and text embeddings for various prompts with regard to the EEG input that is forwarded through the EEG encoder.
+We computed those gradients with regard to the input in the frequency domain for easier analysis.
+Gradients were averaged across the validation set of TUAB.
+Analysis of prompts containing specific frequency descriptions reveals characteristic gradient patterns.
+The prompt "Excessive beta activity" (Figure 10) produces elevated gradient responses in the beta frequency range (14-30 Hz).
+For the more spatially specific prompt "20 Hz beta spindles in central regions" (Figure 11), gradients concentrate in the central electrode locations (C3, Cz, C4) with a peak at approximately 20 Hz, demonstrating the model’s ability to capture both spatial and frequency-specific features mentioned in the text.
+The prompt "Left temporal sharp waves with a frequency of 6 Hz" (Figure 12) produces gradients with lateralized patterns, showing higher magnitudes in left temporal electrodes (T3, T5) with a prominent peak around 6 Hz, indicating successful spatial-frequency alignment with the textual description.
+Figure 13 shows the gradient patterns when the model aligns EEG signals with "This is a normal recording" embeddings.
+The frequency gradient analysis reveals a positive peak around 8-9 Hz in the alpha range, indicating the model identifies healthy recordings with increased alpha band activity.
+This pattern suggests the model has learned to identify normal EEG patterns through alpha frequency features.
+These gradient visualizations provide preliminary evidence that EEG-CLIP learns clinically relevant spatial-temporal patterns rather than purely relying on spurious correlations.
+The distinct gradient patterns between different prompts suggest the model captures meaningful neurophysiological differences.
+However, more sophisticated interpretability methods and validation with clinical experts would be needed to fully understand the clinical relevance of these learned representations.
 5 Discussion
@@
 Most notably, EEG-CLIP demonstrates zero-shot classification capabilities, achieving 0.755 balanced accuracy on pathology detection using only natural language prompts without any task-specific training.
-In low-data regimes, our approach shows substantial benefits over models trained from scratch or pretrained on irrelevant tasks, suggesting efficient capture of generalizable features.
+In low-data regimes, our approach shows substantial benefits over models trained from scratch or pretrained on alternative tasks, suggesting efficient capture of generalizable features.
 Our ablation studies reveal that while using all report sections yields the best overall performance, specific sections provide distinct advantages for certain tasks.
@@
 Despite these promising results, several limitations warrant discussion.
-Our experiments were limited to the TUAB dataset, which may not capture the full diversity of EEG patterns and neurological conditions.
+A significant limitation is our primary reliance on TUAB for evaluation, which raises valid concerns about dataset-specific biases and overfitting.
 Clinical reports in the dataset vary in quality, detail, and structure, potentially limiting the model’s ability to learn precise alignments.
-Additionally, our current approach treats entire EEG recordings and their corresponding reports as aligned pairs, whereas more fine-grained temporal alignment could further improve performance.
+Additionally, our current approach treats entire EEG recordings and their corresponding reports as aligned pairs, whereas more fine-grained temporal alignment between specific EEG segments and relevant report sections could improve performance, as demonstrated by recent Multiple Instance Learning frameworks (Gijsen & Ritter, 2024).
+While we implemented validation strategies including held-out evaluation sets, cross-validation for few-shot experiments, and evaluation across four diverse tasks, broader multi-site validation across different hospital systems, recording protocols, patient populations, and clinical practices would strengthen generalizability claims.
 Future work could explore methods for aligning specific EEG segments with relevant sentences in clinical reports, scaling to larger and more diverse EEG datasets, incorporating additional data modalities, and developing methods to interpret the learned representations in terms of clinically meaningful EEG patterns.
-The development of foundation models for biosignals, analogous to large language models or vision transformers, represents an exciting direction for this research.
+The field would benefit from standardized multi-site evaluation protocols for EEG-language models to better assess generalizability across diverse clinical settings.
 In conclusion, EEG-CLIP demonstrates the feasibility of contrastive learning between EEG signals and natural language descriptions for developing more general and transferable EEG representations.
-While this represents just an initial exploration, the approach opens up new possibilities for leveraging unstructured clinical text to enhance deep learning models for EEG analysis, potentially leading to more flexible, efficient, and powerful tools for neurological assessment and research.
+While this work represents an initial exploration, the approach opens up new possibilities for leveraging unstructured clinical text to enhance deep learning models for EEG analysis, potentially leading to more flexible and data-efficient tools for neurological assessment and research.